# Bounding the Treewidth of Outer $k$-Planar Graphs via Triangulations

### v1

@@ 4.1 Treewidth
We perform the following two modifications of $T$ for each face $f$; see Fig. 5.
@@
  • If $f$ has two children in $T$, assume that piercing number of $\lambda=\{v_{1},v_{3}\}$ is less or equal to that of $\lambda^{\prime}=\{v_{2},v_{3}\}$ and that $\lambda$ is the common link of $f$ and its left child face. If there are $l>0$ bent edges $\{u_{1},w_{1}\},\dots,\{u_{l},w_{l}\}$ w.r.t. $f$, then we subdivide the edge of $T$ between $b_{f}$ and its left child by additional $l$ copies of $b_{f}$. These secondary copy bags are numbered $b_{f,1},\dots,b_{f,l}$ from top to bottom; see Fig. 5.
@@
  • If $e$ is a short edge, that is, it is incident to a vertex of $f$, namely to $v_{1}$ or $v_{2}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ by one bag (as $v_{5}$ and $v_{7}$ in Fig. 6(a)) to the primary copy bag of $b_{f}$.
  • If $e$ is a long edge that pierces the link $\{v_{1},v_{2}\}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ all the way to $b_{f}$; see Fig. 6(b).
  • If $e=\{u_{i},w_{i}\}$ for some $i\in\{1,\dots,l\}$, then we lift both endpoints of $e$. First, we lift the highest occurrence of $u_{i}$ from the left subtree to $b_{f,i}$. Second, we lift the highest occurrence of $w_{i}$ in the right subtree of $b_{f}$ to $b_{f}$. Third, we copy $w_{i}$ into each bag from $b_{f}$ down to $b_{f,i}$; see Fig. 6(c). In this way, the two endpoints of $e$ meet in $b_{f,i}$.
@@
First, we show the bound for secondary copy bags. Recall that, for $i\in\{1,\dots,l\}$, we have added to $b_{f,i}$ from below $i$ endpoints of the bent long edges w.r.t. $f$ and from above $l+1-i$ endpoints of bent long edges; see Fig. 6(c). In addition, we have added a vertex for each lineal long edge. Therefore, the number of added vertices is at most one plus the number of long edges w.r.t. $f$ that pierce the link $\{v_{1},v_{3}\}$. Recall that the link $\{v_{1},v_{3}\}$ was chosen to have at most as many piercing edges as $\{v_{2},v_{3}\}$. Because each inner link of $\mathcal{T}$, including $\{v_{1},v_{2}\}$, is pierced at least once, $\{v_{1},v_{3}\}$ is pierced at most $(c-1)/2$ times. Thus, we have added at most $(c+1)/2$ vertices to $b_{f,i}$, and so $|b_{f,i}|\leq(c+7)/2$.
Now, we bound the number of vertices that are added to the primary (the bottommost) copy bag $b^{\prime}$ of $b_{f}$. Note that such a bag is a parent bag of some original bag $b_{g}$ such that the faces $f$ and $g$ are adjacent. There are two types of vertices that are added to $b^{\prime}$. The first type consists of the endpoints of the lineal long edges of $b_{g}$ (at most $c/2$). The second one is just one vertex $w$ of the face $g$ that is not shared with $f$ (for example, the endpoint $v_{5}$ of $g$ in Fig. 6(a)). We have added $w$ only if there are edges in $E(G)\setminus E(\mathcal{T})$ with lower endpoint $w$ (lower w.r.t. $T$). Each such edge contributes to the piercing number of $g$. Hence, the number of lineal long edges w.r.t. $g$ is at most $(c-1)/2$. In both cases ($w\in b^{\prime}$ and $w\not\in b^{\prime}$), we have added at most $(c+1)/2$ vertices to $b^{\prime}$, so $|b^{\prime}|\leq(c+7)/2$.
@@ 4.2 Separation Number
First, we find a “centroid” triangle of $\mathcal{T}$ as follows. Let $T$ be the tree that is the weak dual of $\mathcal{T}$. It is well known that every tree contains a vertex such that, after removing it, the number of vertices in each subtree is at most half of the original tree. Let $f=\{u,v,w\}$ be the triangle corresponding to such a vertex of $T$. We partition $V\setminus f$ into three disjoint sets $V_{1},V_{2},V_{3}$; see Fig. 7(a). We may assume that $V_{1}$ is the largest among the three sets.
@@
Now, we construct the desired separation with the help of the sets $V_{1},V_{2},V_{3}$. As $\lambda=\{u,v\}$ is a link of $\mathcal{T}$, at most $c$ edges pierce $\lambda$, connecting vertices in $V_{1}$ with vertices in $V_{2}\cup\{w\}\cup V_{3}$. Let $S$ be the set of endpoints of the piercing edges on the latter side. Let $A=V_{1}\cup\{u,v\}\cup S$ and $B=V_{2}\cup\{u,v,w\}\cup V_{3}$; see Fig. 7(b). We claim that $(A,B)$ is a balanced separation of order at most $c+2$.
@@ 5 Lower Bounds
In this section, we complement the results in Section 4, by giving lower bounds for the separation number and treewidth of outer $k$-planar graphs. To show them, we use grid-like graphs called stacked prisms. The $m\times n$ stacked prism $Y_{m,n}$ is the (planar) graph obtained by connecting all pairs vertices in the topmost and bottommost row of the $m\times n$ grid that are in the same column; see Fig. 8(a).
@@
We first observe that, if $m$ is even, $\operatorname{lcr}^{\circ}(Y_{m,n})$ is at most $22n-$ 2. If $m$ is even, we can place the $m$ rows of $Y_{m,n}$, each of length $n$, one after the other, alternating in their direction, around a circle; see Fig. 8(b). The grid has two types of edges: the row edges and the column edges, which connect vertices in the same row or column, respectively. As Fig. 8(b) shows, the row edges have no crossing. For each $i\in[n]$, there is a column edge $e_{i}$ that spans $22i-$ 2 other vertices along the perimeter of the circle. Each of these vertices is incident to exactly one column edge that crosses $e_{i}$. Hence, every column edge has at most $22n-$ 2 crossings, and $\operatorname{lcr}^{\circ}(Y_{m,n})\leq 2n-2$.

### v2

@@ 4.1 Treewidth
We perform the following two modifications of $T$ for each face $f$; see Figure 5.
@@
  • If $f$ has two children in $T$, assume that piercing number of $\lambda=\{v_{1},v_{3}\}$ is less or equal to that of $\lambda^{\prime}=\{v_{2},v_{3}\}$ and that $\lambda$ is the common link of $f$ and its left child face. If there are $l>0$ bent edges $\{u_{1},w_{1}\},\dots,\{u_{l},w_{l}\}$ w.r.t. $f$, then we subdivide the edge of $T$ between $b_{f}$ and its left child by additional $l$ copies of $b_{f}$. These secondary copy bags are numbered $b_{f,1},\dots,b_{f,l}$ from top to bottom; see Figure 5.
@@
  • If $e$ is a short edge, that is, it is incident to a vertex of $f$, namely to $v_{1}$ or $v_{2}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ by one bag (as $v_{5}$ and $v_{7}$ in Figure 6(a)) to the primary copy bag of $b_{f}$.
  • If $e$ is a long edge that pierces the link $\{v_{1},v_{2}\}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ all the way to $b_{f}$; see Figure 6(b).
  • If $e=\{u_{i},w_{i}\}$ for some $i\in\{1,\dots,l\}$, then we lift both endpoints of $e$. First, we lift the highest occurrence of $u_{i}$ from the left subtree to $b_{f,i}$. Second, we lift the highest occurrence of $w_{i}$ in the right subtree of $b_{f}$ to $b_{f}$. Third, we copy $w_{i}$ into each bag from $b_{f}$ down to $b_{f,i}$; see Figure 6(c). In this way, the two endpoints of $e$ meet in $b_{f,i}$.
@@
First, we show the bound for secondary copy bags. Recall that, for $i\in\{1,\dots,l\}$, we have added to $b_{f,i}$ from below $i$ endpoints of the bent long edges w.r.t. $f$ and from above $l+1-i$ endpoints of bent long edges; see Figure 6(c). In addition, we have added a vertex for each lineal long edge. Therefore, the number of added vertices is at most one plus the number of long edges w.r.t. $f$ that pierce the link $\{v_{1},v_{3}\}$. Recall that the link $\{v_{1},v_{3}\}$ was chosen to have at most as many piercing edges as $\{v_{2},v_{3}\}$. Because each inner link of $\mathcal{T}$, including $\{v_{1},v_{2}\}$, is pierced at least once, $\{v_{1},v_{3}\}$ is pierced at most $(c-1)/2$ times. Thus, we have added at most $(c+1)/2$ vertices to $b_{f,i}$, and so $|b_{f,i}|\leq(c+7)/2$.
Now, we bound the number of vertices that are added to the primary (the bottommost) copy bag $b^{\prime}$ of $b_{f}$. Note that such a bag is a parent bag of some original bag $b_{g}$ such that the faces $f$ and $g$ are adjacent. There are two types of vertices that are added to $b^{\prime}$. The first type consists of the endpoints of the lineal long edges of $b_{g}$ (at most $c/2$). The second one is just one vertex $w$ of the face $g$ that is not shared with $f$ (for example, the endpoint $v_{5}$ of $g$ in Figure 6(a)). We have added $w$ only if there are edges in $E(G)\setminus E(\mathcal{T})$ with lower endpoint $w$ (lower w.r.t. $T$). Each such edge contributes to the piercing number of $g$. Hence, the number of lineal long edges w.r.t. $g$ is at most $(c-1)/2$. In both cases ($w\in b^{\prime}$ and $w\not\in b^{\prime}$), we have added at most $(c+1)/2$ vertices to $b^{\prime}$, so $|b^{\prime}|\leq(c+7)/2$.
@@ 4.2 Separation Number
First, we find a “centroid” triangle of $\mathcal{T}$ as follows. Let $T$ be the tree that is the weak dual of $\mathcal{T}$. It is well known that every tree contains a vertex such that, after removing it, the number of vertices in each subtree is at most half of the original tree. Let $f=\{u,v,w\}$ be the triangle corresponding to such a vertex of $T$. We partition $V\setminus f$ into three disjoint sets $V_{1},V_{2},V_{3}$; see Figure 7(a). We may assume that $V_{1}$ is the largest among the three sets.
@@
Now, we construct the desired separation with the help of the sets $V_{1},V_{2},V_{3}$. As $\lambda=\{u,v\}$ is a link of $\mathcal{T}$, at most $c$ edges pierce $\lambda$, connecting vertices in $V_{1}$ with vertices in $V_{2}\cup\{w\}\cup V_{3}$. Let $S$ be the set of endpoints of the piercing edges on the latter side. Let $A=V_{1}\cup\{u,v\}\cup S$ and $B=V_{2}\cup\{u,v,w\}\cup V_{3}$; see Figure 7(b). We claim that $(A,B)$ is a balanced separation of order at most $c+2$.
@@ 5 Lower Bounds
In this section, we complement the results in Section 4, by giving lower bounds for the separation number and treewidth of outer $k$-planar graphs. To show them, we use grid-like graphs called stacked prisms. The $m\times n$ stacked prism $Y_{m,n}$ is the (planar) graph obtained by connecting all pairs vertices in the topmost and bottommost row of the $m\times n$ grid that are in the same column; see Figure 8(a).
@@
We first observe that, if $m$ is even, $\operatorname{lcr}^{\circ}(Y_{m,n})$ is at most $22n-$ 2. If $m$ is even, we can place the $m$ rows of $Y_{m,n}$, each of length $n$, one after the other, alternating in their direction, around a circle; see Figure 8(b). The grid has two types of edges: the row edges and the column edges, which connect vertices in the same row or column, respectively. As Figure 8(b) shows, the row edges have no crossing. For each $i\in[n]$, there is a column edge $e_{i}$ that spans $22i-$ 2 other vertices along the perimeter of the circle. Each of these vertices is incident to exactly one column edge that crosses $e_{i}$. Hence, every column edge has at most $22n-$ 2 crossings, and $\operatorname{lcr}^{\circ}(Y_{m,n})\leq 2n-2$.